OpenStudy (anonymous):

using an addition or subtraction formula sin (43π/12)

OpenStudy (ybarrap):

$$ \sin(A+B)=\sin A\cos B+\cos A\sin B $$ \(\cfrac{43\pi}{12}\) is \(-75^\circ\) in degrees So $$ \sin(-75)=\sin(15-90)=\sin (15)\cos (-90)+\cos (15)\sin (-90)\\ =-\cos (15) $$ \(15^\circ\) is \(\cfrac{\pi}{12}\). $$ \cfrac{\pi}{12}=\cfrac{\pi}{3}-\cfrac{\pi}{4}\\ $$ Using a similar trigonometric identity for the cos sum of angle formula, we finally arrive at our answer: $$ -\cos(A+B)=-\cos A \cos B + \sin A \sin B\\ -\cos \cfrac{1}{12}=-\cos \cfrac{\pi}{3} \cos \cfrac{+\pi}{4} - \sin \cfrac{\pi}{3} \sin \cfrac{-\pi}{4}\\ =-\left (\cfrac{1}{2}\cfrac{\sqrt 2}{2}+\cfrac{\sqrt 3}{2}\cfrac{\sqrt 2}{2}\right )\\ =-\cfrac{1+\sqrt 3}{2\sqrt 2} $$ Make sense?

OpenStudy (anonymous):

yes thank you for being for helping me. Appreciate it